OpenStudy (anonymous):

so this is due tomorrow so please help. Problem- The circumference of a circular region is pie mi. What type of number best describes the diameter of the circle? Explain your answer.

ganeshie8 (ganeshie8):

Notice that the formula for circumference of circle is \(\large \pi D\). Since you're given that the circumference is \(\large \pi\), set them equal to each other solve the diameter \(\large D\) : \[\large \pi D = \pi \\ \large D = ?\]

OpenStudy (anonymous):

I still don't really get it

OpenStudy (triciaal):

circumference = pi mi =pi D where D is the diameter therefore d = mi it seems to me that mi means a complex number

OpenStudy (anonymous):

but there is only a choice of rational, interger,whole number, or irrational

ganeshie8 (ganeshie8):

solve D

OpenStudy (anonymous):

oh its 1